Find the useful power used in pumping 3425m cube of water per hour from a well 8 metres deep to the surface,supposing 40 percent of the horse power during pumping is wasted.What is the horse power of the engine?
1
Expert's answer
2020-03-02T10:20:28-0500

Net power used"=\frac{mgh}{t}" ,where "m" is mass of water uplifted,"h" be height and "t" be the time required.

"t=1 h=3600 sec"

Mass uplifted"=" Density"\times" Volume"=1000\times 3425=3.425\times 10^6 Kg"

Net useful power"=\frac{3.425\times 10^6 \times 9.8\times 8}{3600}=7.458 \times 10^4 W" "=\frac{7.458\times 10^4}{746}=100 hp"

And it is "100-40=60" % of total engine power.

So,"\frac{60}{100}\times P_{net}=100"

"P_{net}=166.67 hp"